EV005

PI: Venturi
Title: 3C75

Session: February 1996 (session 1)
Frequency: 18 cm
Mode: B
Obs time (UT): 46 EU 10:30(15/02)-20:00(15/02)
Stations: SLNJWShUr+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Mon Feb 19 08:26:25 EST 1996) Minor failures
High winds during middle of experiment. Tracking poor from 1030 to 1100UT (up to 6arcmin error in a 30 arcmin FWHM beam). Antenna stowed 1100UT to 1500UT, winds about 15m/s, well above antenna wind limit. Phase-cals switched off, were increasing Tsys significantly in some channels. No other known problems.

Urumqi (Lee Suqin at Wed Feb 21 09:13:07 EST 1996) Success
Cal temprature is 103k and the system temprature is about 210k. It is cloudy. Observation is OK.

Jodrell (Dave Henstock at Thu Feb 22 06:51:40 EST 1996) Minor failures
Bad RFI in observing band (VCs 9,10 also sometimes 11,12 seriously affected). Tsys measurements were made until Day 46 12:40, when was cal diode switched off as it was probable cause of field system crashed. Those Tsys measurements that were made will only be crudely calibrated anyway due to the bad RFI. Note wrong telescope in schedule (should be JODRELL1 not JODRELL2) so beware when correlating. Other problems: Day 46 10:58-12:00 Field System crashed; passes in this time range will be lost 19:43 This scan cancelled by MERLIN as elevation too low. Tsys is in Jy, need to multiply measured Tsys by 1.23 for better calibration.

Cambridge (Dave Henstock at Thu Feb 22 06:54:20 EST 1996) Minor failures
Note that there are no Tsys measurements for Cambridge. Specific Problems: Day 46 12:34-12:47 and 14:19-14:21 Field system crashed, sone/all data between these times will have been lost. 19:43 Scan cancelled by MERLIN as elevation too low. Tsys is 212 Jy (from EVN status).

Medicina (Venturi Tiziana at Thu Feb 22 10:02:56 EST 1996) Minor failures
No problems.

Noto (Corrado Trigilio at Sat Feb 24 06:30:29 EST 1996) Success
No problem reported.

Westerbork (foley at Mon Feb 26 11:47:34 EST 1996) Success
Field system giving false alarms of "Reproduce track B does not check with request"

Shanghai (Huang Xinyong at Fri Mar 1 04:34:20 EST 1996) Success
Observations were fine, weather overcast. Clock offset = -0.9 usec, phase cal signal present.

Calibration (Fredrik T Rantakyro at Thu Sep 26 10:10:48 EDT 1996) Minor failures
An ANTAB formatted file for the experiment can be found on VLBEER